CALL FOR PAPERS Seventh Scandinavian Workshop on Algorithm Theory (SWAT’2000) July 5–7, 2000 Bergen, Norway

                   http://www.ii.uib.no/swat2000

The workshop, which alternates with the Workshop on Algorithms and Data Structures (WADS), is intended as a forum for researchers in the area of design and analysis of algorithms and data structures. We invite submissions of papers presenting original research on algorithms and data structures in all areas, including computational geometry, parallel and distributed computing, graph theory, computational biology, and combinatorics. Contributors should preferably use the electronic submission, see the web page http://www.ii.uib.no/swat2000, or alternatively send five hard copies of a full paper not exceeding 10 pages (plus optionally a clearly marked appendix to be read at the discretion of the program committee) to

Magnús M. Halldórsson Department of Informatics University of Bergen N-5020 Bergen, NORWAY

Important dates:

Submission of Papers  February 15, 2000
Notification of Acceptance March 24, 2000
Final Version due  April 25, 2000

The submissions should give the author’s email address and fax number if available. Proceedings will be published in the Springer-Verlag series Lecture Notes in Computer Science. The final versions of accepted papers must therefore arrive in electronic form, following the LLNCS standard. For more information please check the web pages or send email to telle@ii.uib.no. Program Committee: Amotz Bar-Noy, Tel-Aviv University Luisa Gargano, University of Salerno Jens Gustedt, INRIA Nancy Magnús M. Halldórsson, chair, U. Iceland & U. Bergen Kazuo Iwama, Kyoto University Klaus Jansen, IDSIA Lugano Jan Kratochvíl, Charles University Baruch Schieber, IBM Research Andrzej Lingas, Lund University Jaikumar Radhakrishnan, Tata Institute R. Ravi, Carnegie Mellon University Jörg-Rudiger Sack, Carleton University Hisao Tamaki, Meiji University Jan Arne Telle, University of Bergen Sven Skyum, University of Aarhus Esko Ukkonen, University of Helsinki Gerhard Woeginger, Technical University Graz

Organizing Committee: Pinar Heggernes, University of Bergen Petter Kristiansen, University of Bergen Fredrik Manne, University of Bergen Jan Arne Telle, chair, University of Bergen